# OpenS2S[Uncaptioned image]: Advancing Fully Open-Source End-to-End Empathetic Large Speech Language Model

### v2

@@ Abstract
[Uncaptioned image] Model https://huggingface.co/CASIA-LM/OpenS2S
@@ 3.2 Supervised Fine-tuning
Through this systematic construction process, we obtain an empathetic speech instruction dataset characterized by multi-dimensional emotion annotation, expressive emotional delivery, and diverse speaker profiles. The statistics of the constructed data are presented in Figure 4. In total, we construct 50k English and 50k Chinese speech-to-speech empathetic samples. The input speech includes three types of paralinguistic information tags (i.e. emotion, gender, age)), and the output speech is fixed as a young female voice responding with different emotions.
To retain the model’s general instruction-following capabilities beyond empathetic dialogue, we further incorporate general-purpose data. Specifically, we extract 50k English instructions from Instruct200K [29], translate them into Chinese using Qwen3-32B-Instruct, and select seed audio with neutral emotional labels to convert these instructions into speech. Applying the same speech response generation process as above, we obtain an additional 100k bilingual speech-to-speech instruction pairs.
@@ 4.1 Speech-to-Text Chat
LLaMA-Omni2111We compare with LLaMA-Omni2-7B-Bilingual for its comparable parameter size and bilingual capability. [18] 3.96 3.46 17.36 3.07 39.46 63.79

### v3

@@ Abstract
[Uncaptioned image] OpenS2S https://huggingface.co/CASIA-LM/OpenS2S
OpenS2S_V1.5111 [October 2025] We have released OpenS2S_V1.5, which demonstrates significant improvements in performance across multiple metrics and practical experiences compared to the original version. https://huggingface.co/CASIA-LM/OpenS2S_V1.5
@@ 3.2 Supervised Fine-tuning
Through this systematic construction process, we obtain an empathetic speech instruction dataset characterized by multi-dimensional emotion annotation, expressive emotional delivery, and diverse speaker profiles. The statistics of the constructed data are presented in Figure 4. In total, we construct 50k English and 50k Chinese speech-to-speech empathetic samples. The input speech includes three types of paralinguistic information tags (i.e. emotion, gender, age), and the output speech is fixed as a young female voice responding with different emotions. [UPDATE (October 2025) In the V1.5 version, we expand the training data using the same approach, introducing approximately 200K English and 200K Chinese speech-to-speech empathetic samples.]
To retain the model’s general instruction-following capabilities beyond empathetic dialogue, we further incorporate general-purpose data. Specifically, we extract 50k English instructions from Instruct200K [29], translate them into Chinese using Qwen3-32B-Instruct, and select seed audio with neutral emotional labels to convert these instructions into speech. Applying the same speech response generation process as above, we obtain an additional 100k bilingual speech-to-speech instruction pairs.[UPDATE (October 2025) In the V1.5 version, we incorporate all data from Instruct200K, resulting in 400K bilingual speech-to-speech instruction pairs.]
@@ 4.1 Speech-to-Text Chat
LLaMA-Omni2222We compare with LLaMA-Omni2-7B-Bilingual for its comparable parameter size and bilingual capability. [18] 3.96 3.46 17.36 3.07 39.46 63.79
@@
UPDATE (October 2025) We test the newly released OpenS2S_V1.5, which is based on OpenS2S and continues the original method for further fine-tuning. This model only extend the data scale during the Empathetic Speech Instruction Tuning phase. The results show that OpenS2S_V1.5 shows significant improvements across most metrics, with two metrics surpassing all baseline models. The results after scaling the training data further validate the effectiveness of the data generation method we proposed.